OpenStudy (anonymous):

A ball is thrown 125 m upwards and then falls back to earth. Neglecting air resistance, how long will it be in the air? The acceleration of gravity is 9.8m/s^2

OpenStudy (anonymous):

do you know your kinematic equations?

OpenStudy (jt950):

S=125 u=0 g=9.8m/s^2 S=ut+(1/2)at^2 125=0+(1/2)(9.8)t^2 t=5.05second

OpenStudy (anonymous):

I thought it was around 5 seconds but it was incorrect. And I do not know the equation

OpenStudy (jt950):

t=5.05second is time for object fall downward, for total time in air = 5.05x2 = 10.1second

OpenStudy (anonymous):

Thank you :)